# Symmetry breaking effects in pion couplings to constituent quarks currents

### v1

@@ 4 Numerical Results
Given that there is an intrinsic ambiguity in defining values for the quark effective masses, results will be shown for different choices of $M_{u},M_{d}$ (and also $M_{s}$). These masses, for the chosen gluon effective propagator, determine $G_{ij}^{\xi}$ (both $G_{ps}$ and $G_{A}$) numerically, and the different sets of parameters (s.o.p.) are shown in Table (1). Sets of parameters $A_{1}$ and $A_{2}$ were extracted from [21]. The third $A_{3}$ is based in often values found in the literature.
@@
In Table (2) values of pseudoscalar coupling constants $G_{ps}$, defined in different kinematical points (off shell or on shell constituent quark and pion) are shown for the different sets of parameters presented in Table (1). Besides the charged pion coupling constant, the neutral pion coupling constant is exhibited with and without flavor states mixings, respectively $G^{ps}_{\pi^{0}}(0)$ and $G^{ps}_{\pi^{0}}(ij)$ from Eqs. (11) and (17). As shown above the neutral pion coupling can be decomposed in the couplings of states $\bar{u}u$ and $\bar{d}d$ states, that might also receive contributions from flavor state mixings - and these are shown in the middle part of this Table. In the last two columns the pseudoscalar coupling constant of the neutral pion coupling to strange quark current $G_{mix,ss}^{ps}$ is exhibited with a parameter $x_{ps}^{m}$ that will be defined below.
Results for s.o.p. $A_{2}$ ($A_{1}$), with the lower (larger) values of quark masses, present the smallest (largest) variation of values for the different kinematical points. The charged pion coupling constants are smaller than the neutral pion ones except for the case of the fourth kinematical point: off shell quarks and on shell pions $(K^{2}=0,Q^{2}=M_{\pi}^{2})$. The neutral pion coupling to the $\bar{u}u$ quark current has larger strength than the coupling to $\bar{d}d$ for all kinematical points, and for both cases, without and with mixings, $G^{ps}_{uu}(0)$ and $G^{ps}_{uu}(ij)$ respectively.
Recent analysis of the pion-nucleon pseudoscalar coupling constant calculated with different approaches [6, 5] exhibit a very similar trend with small numerical difference between charged and neutral pion couplings to nucleons. The neutral pion -nucleon coupling constants are slightly larger than the charged pion coupling coupling constant. For the sake of comparison of the behavior, the following values can be extracted from the calculation with the EFT approach [6]: $g_{\pi^{0}pp}=13.22688$, $g_{\pi^{0}nn}=13.39731$, $g_{\pi^{\pm}np}=13.22611$. According to the quark model, proton and neutron structures suggest that the corresponding coupling to the pion might be a sort of sum or average for the pion coupling to corresponding states, $p\sim uud$ and $n\sim ddu$ respectively. However, due to the choice of the renormalization condition ($G^{ps}_{\pi^{p}m}=13$), the resulting neutral pion coupling to nucleons, as a first simplified approach, might be considered as a simple average:
@@
Concerning the charged pion coupling, the behavior shown in the Table is consistent with the results for the charged pion-neutron-proton and neutral pion-proton coupling constants such that that is $g_{\pi^{\pm}np}<g_{\pi^{0}pp},g_{\pi^{0}nn}$ and $g_{\pi^{\pm}np}\sim g_{\pi^{0}pp}$ [6, 5]. Along these lines, it is also expected that $g_{\pi^{0}nn}>g_{\pi^{0}pp}$. This hierarchy in pion-nucleon coupling constants, however, does not happen with the values shown in the Table for (22) for the model of (22), since it would require $G_{\pi^{0}dd}>G_{\pi^{0}uu}$ i.e. the opposite behavior to the one shown in the Table. It is shown below that mixings can mend this behavior.
By varying the strength of the mixing interactions by a constant uniform factor $x^{m}_{ps}$, the pseudoscalar coupling constants $G^{ps}_{uu}$ and $G^{ps}_{dd}$ can flip their hierarchy if $x_{ps}^{m}<0$. Therefore mixing parameters were redefined by:
@@
$(M_{f}^{2},0)$ 1.1602 1.1601 1.1601 1.1577 1.1536 1.1626 1.1666 0.0001
$(M_{f}^{2},M_{\pi}^{2})$ 1.3647 1.3547 1.3547 1.3517 1.3469 1.3578 1.3626 0.0002
$(0,0)$ 0.9030 0.9030 0.9029 0.9052 0.9020 0.9008 0.9040 -0.0001
$(0,M_{\pi}^{2})$ 0.9296 0.9278 0.9278 0.9305 0.9273 0.9251 0.9283 -0.0001
As noted before, the positive pion absorption (negative pion emission) coupling constant in both channels, $G_{du}^{\xi}(K^{2},Q^{2})$ ($\xi=ps,A$), might be different from the negative pion absorption (positive pion emission) coupling constant, $G_{ud}^{\xi}(K^{2},Q^{2})$ in the case of non-degenerate quark masses. The difference between these two charged pion coupling constants, for both pseudoscalar and axial couplings, is shown in Fig. (2), by means of the following (on shell quarks) ratios:
@@
Two cases are exhibited, $Q^{2}=M_{\pi^{\pm}}^{2}$ and $Q^{2}=0$. It is seen that the variation is very small and basically linear, i.e. goes with the difference of masses $M_{d}-M_{u}$ in the leading order. The relative variation of the axial coupling constant is smaller than the pseudoscalar. The case of on shell or massive (off shell or massless) pion presents a larger variation for the pseudoscalar (axial) coupling constant.
@@ 5 Summary and conclusions
Pseudoscalar and axial charged and neutral pion -constituent quark coupling constants were derived and investigated by considering quark mass non-degeneracy within an approach proposed previously by the author. These coupling constants were calculated for different sets of parameters of constituent quark masses and in different kinematical points: for on shell or off shell pions and constituent quarks. The pseudoscalar coupling constant calculated for off shell (or massless) charged pion and on shell constituent quark was fixed as a renormalization condition by identifying it to a phenomenological value $G^{ps}_{\pi^{\pm}}(K^{2}=M_{f}^{2},Q^{2}=0)=13$. With this condition, the resulting axial coupling constant is obtained with values close to the value adopted for the CQM, i.e. $g_{A}\sim 1$. Being all the coupling constants functions of the quark constituent masses (and not the small current quark masses), the Goldberger-Treiman relation at the quark level presents, in fact, a large correction for the values $G_{ps}=13$ and $g_{A}=1$. Results suggest that lower values of quark effective masses lead to smaller differences between the neutral and charged pion couplings (both axial and pseudoscalar). A very small difference between charged and neutral pion couplings has been noted in the pion-nucleon pseudoscalar coupling constant. The neutral pion coupling to quark currents $\bar{u}u$ and $\bar{d}d$ was compared to neutral pion coupling to protons and neutrons (respectively) and this comparison suggests mixing interactions stronger than those provided by flavor symmetry breaking are needed. This could be provided by instanton induced ’t Hooft-like interactions.
@@
The resulting charged pion coupling constants from the quark determinant present full U(1) and CP invariance with $G_{11}=G_{22}$. It corresponds to an average of emitted and absorbed (positive or negative) pion, being therefore the same for each charge eigenstate or for each type of process (emission or absorption). For non-degenerate quark masses, there appears a small difference between absorption and emission of a positive or negative pion. This difference is proportional - in the leading order - to the quark mass difference $(M_{d}-M_{u})$. However, high precision experimental results/fittings would be needed to identify such difference.
@@
Interactions that mix neutral flavor eigenstates $i,j=0,3,8$, leading to usual $\pi^{0}-\eta-\eta^{\prime}$ mixing, have been taken into account for both the quark currents and for the neutral pion. The shifts in the neutral pion coupling to the quark current $\bar{u}u$ are opposite to the shifts in the coupling to the current $\bar{d}d$. Besides that, a quite weak neutral pion coupling to the strange current $\bar{s}s$ was obtained for both axial and pseudoscalar cases. The relative strength of these couplings to strangeness, with respect to the pseudoscalar and axial ones, is very small $\sim 10^{-3},10^{-4}$, so that their contributions for neutral pion couplings to strange baryons would receive very tiny shifts only possibly identified with high precision measurements. Although the hierarchy of the strength of the flavor dependent coupling constants might nearly follow the one from the nucleon-pion couplings, the behavior of these coupling constants by inserting these couplings into the three-quark bound system for the nucleon can bring further insights into this subject. Besides that, at finite temperature, baryon density or external electromagnetic fields, all these couplings must receive non-negligible contributions. This type of investigation might seem not appropriated for the understanding of hadron interactions because quarks are confined in hadrons. However it is important to emphasize that confinement operates in color degrees of freedom whereas flavor (isospin) cannot be (completely) confined, otherwise all the quark-based description of hadrons, and maybe QCD, would be invalidated.
@@ Acknowledgements
The author thanks a F.L.B. is a member of INCT-FNA, Proc. 464898/2014-5 and he acknowledges partial support from CNPq-312750/2021 and CNPq-407162/2023-2.

### v2

@@ 4 Numerical Results
Given that there is an intrinsic ambiguity in defining values for the quark effective masses, results will be shown for different choices of $M_{u},M_{d}$ (and also $M_{s}$). These masses, for the chosen gluon effective propagator, determine $G_{ij}^{\xi}$ (both $G^{ps}$ and $G^{A}$) numerically, and the different sets of parameters (s.o.p.) are shown in Table (1). In the present work, we consider quark effective masses - sets of parameters $A_{1}$ and $A_{2}$ obtained from a (flavor-dependent) NJL model as presented in [24]. In NJL-type models there is need to fix chosen observables/quantities to be mass independent and this choice may bring differences in the results, for example in Refs. [38], being the neutral and charged pion masses chosen to fit the model parameters. In the present calculation however, the renormalization condition presented in Eq. (19) is a further "constraint" for the resulting effective action and numerical results. The third s.o.p., $A_{3}$, is based in values often found in the literature.
@@
In Table (3) values of pseudoscalar coupling constants $G_{ps}$, defined in different kinematical points (off shell or on shell constituent quark and pion) are shown for the different sets of parameters presented in Table (1). The following kinematical points have been considered:
$S_{1}:(K^{2},Q^{2})=(M_{f}^{2},0),$ $\;\;\;S_{2}:(K^{2},Q^{2})=(M_{f}^{2},M_{\pi}^{2}),$ (22)
$S_{3}:(K^{2},Q^{2})=(0,0),$ $\;\;\;S_{4}:(K^{2},Q^{2})=(0,M_{\pi}^{2}).$
Besides the charged pion coupling constant, the neutral pion coupling constant is exhibitted with and without flavor states mixings, respectively $G^{ps}_{\pi^{0}}(0)$ and $G^{ps}_{\pi^{0}}(ij)$ from Eqs. (11) and (17). As shown above the neutral pion coupling can be decomposed in the couplings of states $\bar{u}u$ and $\bar{d}d$ states, that might also receive contributions from flavor state mixings - and these are shown in the middle part of this Table. In the last two columns the pseudoscalar coupling constant of the neutral pion coupling to strange quark current $G_{mix,ss}^{ps}$ is exhibitted with a parameter $x_{ps}^{m}$ that will be defined below.
Results for s.o.p. $A_{2}$ ($A_{1}$), with the lower (larger) values of quark masses, present the smallest (largest) variation of values for the different kinetical points. The charged pion coupling constants are smaller than the neutral pion ones except for the case of the fourth kinematical point: off shell quarks and on shell pions $(K^{2}=0,Q^{2}=M_{\pi}^{2})$. The neutral pion coupling to the $\bar{u}u$ quark current has larger strength than the coupling to $\bar{d}d$ for all kinematical points, and for both cases, without and with mixings, $G^{ps}_{uu}(0)$ and $G^{ps}_{uu}(ij)$ respectively.
Recent analysis were done of the pion-nucleon pseudoscalar coupling constant with different approaches [6, 7]. With different definitions of coupling constants, usually denoted by $f_{\pi^{a}NN^{\prime}}$ for $a=0,\pm$ and $N,N^{\prime}=n,p$, results exhibit a somehow similar trend with small numerical difference between charged and neutral pion couplings to nucleons. For the sake of comparison of the behavior, some values are reproduced in Table (2). The first one is the value extracted from Ref. [7], the second and third lines correspond to values presenting smallest standard deviation confidence level from Ref. [6]. In spite of the differences in absolute values the analysis will be focused on the relative values. Whereas the coupling constant of neutral pion to neutrons is larger than the coupling constant to protons in both approaches, there is a considerable difference between the results of these approaches concerning the charged pion coupling to the nucleon.
Table 2: Neutral and charged pion-nucleon pseudoscalar coupling constants extracted from different recent works [6, 7].
$G_{\pi^{0}nn}$ $G_{\pi^{0}pp}$ $G_{\pi^{\pm}np}$
$1$- [7] 13.397 13.227 13.226
$2$- [6] 15.455 12.511 18.571
$3$- [6] 13.452 13.174 18.571
According to most versions of the quark model, proton and neutron couple to mesons by means of quarks in an addictive way [15, 35]. Along this idea, the pion coupling to nucleons might also arise in a sort of average for the pion coupling to corresponding quark currents, $p\sim uud$ and $n\sim ddu$. Due to the choice of the renormalization condition ($G^{ps}_{\pi^{\pm}}=13$), the resulting neutral pion coupling to nucleons, as a first simplified approach, might be considered as a simple average:
@@
The behavior shown in the Table (3) is partially consistent with the results for the charged pion-neutron-proton and neutral pion-neutron coupling constants of Ref. [6] by considering these averages, such that that is $g_{\pi^{\pm}np}>g_{\pi^{0}nn}$. However, the behavior of values in Table (3) is rather similar to [7] when comparing charged pion to neutral-pion proton coupling constants, i.e. $g_{\pi^{\pm}np}\sim g_{\pi^{0}pp}$ and $g_{\pi^{\pm}np}<g_{\pi^{0}pp}$. For the results shown in the Table (3) it would be expected that $g_{\pi^{0}nn}<g_{\pi^{0}pp}$ and this is not compatible with the results shown in Table (2). Overall, there is only partially agreement between results of these two quoted works and the present one. It is shown below that mixings can modify the relative behavior of $G_{uu}^{ps}(ij)$ and $G_{dd}^{ps}(ij)$, and then $g_{\pi^{0}nn}$ and $g_{\pi^{0}pp}$. NJL-type models, have been considered in [39] where an hierarchy of coupling constant values has been obtained that is (at least partially) in disagreement with all these quoted works and with the present one.
By varying the strength of the mixing interactions by an arbitrary constant uniform factor $x^{m}_{ps}$, the pseudoscalar coupling constants $G^{ps}_{uu}$ and $G^{ps}_{dd}$ can flip their hierarchy (suitable for making $g_{\pi^{0}pp}<g_{\pi^{0}nn}$) if $x_{ps}^{m}<0$. Therefore mixing parameters were redefined by:
@@
$S_{1}$ 1.160(1) 1.160(2) 1.160(2) 1.158(1) 1.154(1) 1.163(1) 1.167(1) 0.001(1)
$S_{2}$ 1.365(1) 1.355(2) 1.355(2) 1.352(1) 1.347(1) 1.358(1) 1.363(1) 0.001(1)
$S_{3}$ 0.903(1) 0.903(2) 0.903(2) 0.905(1) 0.902(1) 0.901(1) 0.904(1) -0.001(1)
$S_{4}$ 0.930(1) 0.928(2) 0.928(2) 0.931(1) 0.927(1) 0.925(1) 0.928(1) -0.001(1)
As noted before, the positive pion absorption (negative pion emission) coupling constant in both channels, $G_{du}^{\xi}(K^{2},Q^{2})$ ($\xi=ps,A$), might be different from the negative pion absorption (positive pion emission) coupling constant, $G_{ud}^{\xi}(K^{2},Q^{2})$ in the case of nondegenerate quark masses. The difference between these two charged pion coupling constants, for both pseudoscalar and axial couplings, is shown in Fig. (2), by means of the following (on shell quarks) ratios:
@@
Two cases are exhibitted, $Q^{2}=M_{\pi^{\pm}}^{2}$ and $Q^{2}=0$. It is seen that the variation is very small and basically linear, i.e. goes with the difference of masses $M_{d}-M_{u}$ in the leading order. The relative variation of the axial coupling constant is smaller than the pseudoscalar. The case of on shell or massive (off shell or massless) pion presents a larger variation for the pseudoscalar (axial) coupling constant.
@@ 5 Summary and conclusions
Pseudoscalar and axial charged and neutral pion -constituent quark coupling constants were derived and investigated by considering quark mass non-degeneracy within a dynamical approach proposed previously by the author. In this approach, coupling constants and form factors are obtained from a quark determinant in the presence of meson fields and quark currents. Gluon effects are considered by means of a gluon effective propagator that dresses background quark currents leading to constituent quark currents. These coupling constants were calculated numerically for different sets of parameters of constituent quark masses and in different kinematical points: for on shell or off shell pions and constituent quarks. The pseudoscalar coupling constant calculated for off shell (or massless) charged pion and on shell constituent quark was fixed as a renormalization condition by identifying it to a phenomenological value $G^{ps}_{\pi^{\pm}}(K^{2}=M_{f}^{2},Q^{2}=0)=13$. With this condition, the resulting axial coupling constant is obtained with values close to the value adopted for the CQM, i.e. $g_{A}\sim 1$. Being all the coupling constants functions of the quark constituent masses (and not the small current quark masses), the Goldberger Treiman relation - as usually defined - at the quark level may present, in fact, a large correction for the values $G_{ps}=13$ and $g_{A}=1$. The differences between charged and neutral pion couplings to the nucleon recently obtained in different approaches are small and this trend is reproduced with the constituent quark couplings in this work. The specific hierarchies of the different coupling constants, however, are different among each other, from [6, 7, 39] and also the present. Results suggest that lower values of quark effective masses lead to smaller differences between the neutral and charged pion couplings (both axial and pseudoscalar). The average of neutral pion couplings to quark currents $\bar{u}u$ and $\bar{d}d$ were compared to neutral pion coupling to protons and neutrons and this comparison suggests mixing interactions should be stronger than those provided by flavor symmetry breaking to reproduce values from other works such as Refs.[6, 7]. This extra mixing could be provided by instanton induced ’t Hooft-like interactions. A more extensive investigation of the role of the specific gluon propagator will be provided elsehwere.
@@
The resulting charged pion coupling constants from the quark determinant present full U(1) and CP invariances with $G_{11}=G_{22}$. It corresponds to an average of emitted and absorbed (positive or negative) pion, being therefore the same for each charge eigenstate or for each type of process (emission or absorption). For non-degenerate quark masses, there appears a small difference between absorption and emission of a positive or negative pion. This difference is proportional - in the leading order - to the quark mass difference $(M_{d}-M_{u})$. However, high precision experimental results/fittings would be needed to identify such difference.
@@
FSB induced interactions, that mix neutral flavor eigenstates, $i,j=0,3,8$, leading to usual $\pi^{0}-\eta-\eta^{\prime}$ mixing, have been taken into account for both the quark currents and for the neutral pion. The shifts in the neutral pion coupling to the quark current $\bar{u}u$ are opposite to the shifts in the coupling to the current $\bar{d}d$. Besides that, a quite weak neutral pion coupling to the strange current $\bar{s}s$ was obtained for both axial and pseudoscalar cases. The relative strength of these couplings to strange current with respect to the pseudoscalar and axial ones, is very small $\sim 10^{-3},10^{-4}$, so that their contributions for neutral pion couplings to strange baryons would receive very tiny shifts only possibly identified with high precision measurements. However, note that this coupling is proportional to the mixing interaction $G_{i\neq j}$ that must receive contributions from ’t Hooft-type interaction and sixth order interactions and these might increase considerably its value. Nevertheless, given that hadrons are strongly interacting systems, small contributions might, possibly, lead to sizeable effects. Although the strength of the flavor dependent constituent quark coupling constants might follow an hierarchy the behavior of these coupling constants by inserting these couplings into the three-quark bound system for the nucleon can bring further insights into this subject. Besides that, at finite temperature, baryon density or external electromagnetic fields, all these couplings must receive non-negligeable contributions. This type of investigation might seem not appropriated for the understanding of hadron interactions because quarks are confined in hadrons. However it is important to emphasize that confinement operates in color degrees of freedom whereas flavor-isospin cannot be (completely) confined, otherwise all the quark-based description of hadrons, and maybe QCD, would be invalidated.
@@ Acknowledgements
The author thanks short discussion with C.D. Roberts in early stages of this work. F.L.B. is a member of INCT-FNA, Proc. 464898/2014-5 and he acknowledges partial support from CNPq-312750/2021-8 and CNPq-407162/2023-2.